A woman is currently 21 years older than her daughter. Eleven years before the current time. the daughter was two-thirds as old as her mother. what is the current age of the mother?
Found 3 solutions by Alan3354, ikleyn, greenestamps:
Answer by Alan3354(69443)

"currently" is superfluous, serves no purpose.
------
Same problem, different numbers.
===================================
Carmen is 12 years old ??? than David. Five years ago the sum of their ages was 28. How old are they now?
------------------
Carmen is 12 years older than David. Five years ago the sum of their ages was 28. How old are they now?
---
The sum of their ages now is 38.
---
C + A = 38
C - A = 12
-------------------- Add

Answer by ikleyn(52803)

.

It is a  TOTAL  and  ABSOLUTE  disharmony with  COMMON  SENSE  to read what you write in your post

                " . . . A woman is currently 21 years older than her daughter."


            +----------------------------------------------------------------------------------+
                In such "age problems", it is assumed (dictated by the common sense)
                that if one person is "N years" older than the other person,
                then it is TRUE forever until these persons are alive.
            +---------------------------------------------------------------------------------+


How you write your post,  it shows  WITHOUT  ANY WORDS  that you  (or the  "sourse"  of this problem)
are unfamiliar with the subject  ("age word problems",  in this case).



Answer by greenestamps(13200)



Eleven years ago, the daughter was 2/3 the age of the mother. So

let 3x = age of mother 11 years ago
then 2x = age of daughter 11 years ago

The difference in their ages 11 years ago (and always!) is 21 years. But the difference between 3x and 2x is x; so x=21.

That means 11 years ago the mother's age was 3x=3(21)=63 and the daughter's age was 2x = 2(21) = 42.

And that means their current ages are 63+11=74 and 42+11=53.

ANSWER: The mother is now 74 years old.